OpenStudy (anonymous):

can someone help me find the antiderivative of e^(4x^(2))?

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

What is the whole problem?

OpenStudy (anonymous):

well i tried substitution but it got a little complicated with the integral. what i needed to do was find the integral of \[\pi \int\limits_{0}^{1} e ^{4}- e ^{4x ^{2}}dx\] and i know what to do with everything but finding the aniderivative of \[e ^{4x ^{2}}\] was getting diffucult

OpenStudy (anonymous):

let: \[u = x^2\] \[\sqrt{u} = x\] du = 2xdx du/x = 2dx \[du/\sqrt{u} = 2dx\] so now: \[\int\limits_{}^{} e ^{x^2} dx = (1/2)\int\limits_{}^{} (e^u)(1/\sqrt{u}) du\] and then i got stuck :(

OpenStudy (anonymous):

why does u=\[x^{2}\] and not 4\[x ^{2}\]?

OpenStudy (amistre64):

e^(x^2) has no elementary function to find

OpenStudy (anonymous):

\[4x ^{2}\]

OpenStudy (amistre64):

unless you got something to counteract that 4x^2 its not doable with what youve been taguth

OpenStudy (amistre64):

taught .. lol, at least i got the right keys

OpenStudy (anonymous):

do you all want the very very original problem? and np lol

OpenStudy (anonymous):

please

OpenStudy (amistre64):

if ther eis an original then yes

OpenStudy (anonymous):

okay hold on it's a little long. thank you all!

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

Sketch the region R bounded by the graphs of the given equations. After showing a typical cross-sectional, find the volume of the solid generated by revolving R about the x-axis. \[y=e ^{2x}\] x=0 \[y=e ^{2}\]

OpenStudy (amistre64):

ugh, this one all over again

OpenStudy (amistre64):

or something close to it lol

OpenStudy (anonymous):

and after that i did \[\pi \int\limits_{0}^{1} r _{1}^{2} - r _{2}^{2}dx\]

OpenStudy (amistre64):

\[(e^{2x})^2=e^{2x}*e^{2x}=e^{2x+2x}=e^{4x }\]

OpenStudy (anonymous):

then pluged it in to get \[\pi \int\limits_{0}^{1} (e ^{2})^{2} - (e ^{2x})^{2}dx\]

OpenStudy (anonymous):

yeah and then i got stuck at finding the antiderivative of \[e ^{4x ^{2}}\]

OpenStudy (amistre64):

then quit trying to find the antiderivative of a function you dont even need

OpenStudy (anonymous):

to find the integral im suppose to

OpenStudy (amistre64):

look up like 5 posts

OpenStudy (amistre64):

see the difference between what your doing and what your spose to do?

OpenStudy (anonymous):

so when i plugged it in it's suppose to be multiplied by 2 and not squared?

OpenStudy (amistre64):

when you power up exponents; the result is they multiply together

OpenStudy (amistre64):

\[(b^n)^p=b^{np}\]

OpenStudy (anonymous):

LOL oh man: \[r = e ^{2x}\] then: \[r^2 = e ^{4x}\] THAT'S A LOT EASIER TO INTEGRATE

OpenStudy (anonymous):

oh okay no wonder!!! >.< thank you thank you both!!

OpenStudy (amistre64):

youre welcome :)